What is the ordered pair using the method of substitution to solve the system of liner equations x-y=3 5x-5y=15
Answer by richwmiller(17219)

They are the same line.
Multiply the first by 5 and you will get the second.
x=y+3
5*(y+3)-5y=15
5y+15-5y=15
5y=5y
y=y